# Quadrature rules with few nodes supported on algebraic curves

### v1

@@ 1. Introduction
for all polynomials $p$ of degree at most $s$. The points in $N$ are called the nodes of the quadrature rule. While one could in principle allow signed weights, positivity ensures monotonicity of the integral, improves numerical stability, and often yields sharper error estimates [13, Conclusion 3.19].
The existence of such rules for arbitrary measures goes back to classical results of Richter [18], Tchakaloff [20], and Bayer–Teichmann [2]. Using Carathéodory’s theorem it follws from these results that a quadrature rule of strength $s$ always exists with at most
@@
    nodes, and a quadrature rule of strength $2s$ for $\mu$ with at most
    $N\leq Ds+p+1$
    nodes.
@@
The question of the minimal number of nodes for general algebraic curves in the plane has been studied by Riener and Schweighofer [18]. Their work recasts the construction of quadrature rules as a conic optimization problem to obtaine refined bounds for measures supported on plane algebraic curves. Subsequent work by Zalar Zalar [21] further sharpened these bounds for certain families of curves using moment matrix methods.
@@
Further connections between moment cones, Hilbert function methods, and quadrature node bounds have been developed by Baldi, Blekherman, and Sinn [1]. A broader geometric perspective on cubature formulas for spheres, quadrics, and polynomially defined domains is provided by Lasserre [15, 16], Sommariva–Vianello [19], and Gustafsson [12].
@@ 3. Quadrature on plane curves
where $R>0$ is chosen so that the feasible set is non-empty and compact. Existence of a such an $R$ follows from existence of quadrature rules per se [17].
@@ 4. Quadrature on rational curves
Since $\mu^{\prime}$ is nondegenerate in degree $sd$ and Gaussian quadrature is optimal for nondegenerate measures (see [18, Thm 9.2])   a quadrature rule exact for degrees up to $m$ requires at least $\lceil(m+1)/2\rceil$ nodes. With $m=sd$, this gives $k\geq\lceil(sd+1)/2\rceil$. Equivalently, if $sd=2\ell-1$ then $k\geq\ell$, and if $sd=2\ell$ then $k\geq\ell+1$. ∎
@@
Again we observe the Lagrange conditions: The Lagrangian is
@@
Thus among the positive roots of $H$, only those with non-positive derivative are candidates for local minima; at most half of the roots can satisfy this sign condition.
@@
Hence any $H(t)$ constructed as above has at most
@@
Among these roots, at most half satisfy $H^{\prime}(t_{i})\leq 0$. Thus the number of local minima is at most
@@
Several natural questions remain open: Firstly, while our estimates are tight in some cases (e.g., lines or certain low-degree curves), it is unclear whether they are optimal for all degrees and strengths. In particular, it seems plausible that the additional counting for places at infinity can be significantly reduced, possibly to a constant independent of $t$, or even to a single point. Secondly, it would seems hopefuly, that extending the present approach to algebraic surfaces or general real varieties of dimension $\geq 2$ may reveal new phenomena. This will likely require refined tools from real algebraic geometry and moment theory to control both local and global contributions to node counts. Thirdly, it seems natural to ask if the set of quadrature rules of a given strength on a fixed curve $\mathcal{C}$ can be represented similarly to the Gaussian case, as was studied in [3].
Finally, our bounds depend only on degree and parametrization data. Incorporating finer invariants, such as singularity structure, genus, or real topology, might lead to sharper or even exact formulas for minimal node counts in broad generality.
These questions point toward a interesting interplay between real algebraic geometry, optimization, and numerical analysis in the study of quadrature rules on algebraic sets.

### v2

@@ 1. Introduction
for all polynomials $p$ of degree at most $s$. The points in $N$ are called the nodes of the quadrature rule. While one could in principle allow signed weights, positivity ensures monotonicity of the integral, improves numerical stability, and often yields sharper error estimates [16, Conclusion 3.19].
The existence of such rules for arbitrary measures goes back to classical results of Richter [21], Tchakaloff [23], and Bayer–Teichmann [2]. Using Carathéodory’s Theorem it follows from these results that a quadrature rule of strength $s$ always exists with at most
@@
    nodes, and a quadrature rule of strength $2s$ for $\mu$ with at most NDs+p+1 nodes.
@@
The question of the minimal number of nodes for general algebraic curves in the plane has been studied by Riener and Schweighofer [21]. Their work recasts the construction of quadrature rules as a conic optimization problem to obtain refined bounds for measures supported on plane algebraic curves. Subsequent work by Zalar [24] further sharpened these bounds for certain families of curves using moment matrix methods.
@@
Further connections between moment cones, Hilbert function methods, and quadrature node bounds have been developed by Baldi, Blekherman, and Sinn [1]. A broader geometric perspective on cubature formulas for spheres, quadrics, and polynomially defined domains is provided by Cools [5, 6], Cools-Rabinowitz [7], Lasserre [18, 19], Sommariva–Vianello [22], and Gustafsson [15].
@@ 3. Quadrature on plane curves
where $R>0$ is chosen so that the feasible set is non-empty and compact. Existence of such an $R$ follows from the existence of quadrature rules per se [20].
@@ 4. Quadrature on rational curves
Since $\mu^{\prime}$ is nondegenerate in degree $sd$ and Gaussian quadrature is optimal for nondegenerate measures (see [21, Thm 9.2])   a quadrature rule exact for degrees up to $m$ requires at least $\lceil(m+1)/2\rceil$ nodes. With $m=sd$, this gives $k\geq\lceil(sd+1)/2\rceil$. Equivalently, if $sd=2\ell-1$ then $k\geq\ell$, and if $sd=2\ell$ then $k\geq\ell+1$. ∎
@@
Again, we observe the Lagrange conditions: The Lagrangian is
@@
Thus, among the positive roots of $H$, only those with non-positive derivative are candidates for local minima; at most half of the roots can satisfy this sign condition.
@@
Hence, any $H(t)$ constructed as above has at most
@@
Among these roots, at most half satisfy $H^{\prime}(t_{i})\leq 0$. Thus, the number of local minima is at most
@@
Several natural questions remain open: Firstly, while our estimates are tight in some cases (e.g., lines or certain low-degree curves), it is unclear whether they are optimal for all degrees and strengths. In particular, it seems plausible that the additional counting for places at infinity can be significantly reduced, possibly to a constant independent of $t$, or even to a single point. Secondly, it would seem hopeful that extending the present approach to algebraic surfaces or general real varieties of dimension $\geq 2$ may reveal new phenomena. This will likely require refined tools from real algebraic geometry and moment theory to control both local and global contributions to node counts. Thirdly, it seems natural to ask if the set of quadrature rules of a given strength on a fixed curve $\mathcal{C}$ can be represented similarly to the Gaussian case, as was studied in [3].
Finally, our bounds depend only on the degree and parametrization data. Incorporating finer invariants, such as singularity structure, genus, or real topology, might lead to sharper or even exact formulas for minimal node counts in broad generality.
These questions point toward an interesting interplay between real algebraic geometry, optimization, and numerical analysis in the study of quadrature rules on algebraic sets.